# A complete proof of the geometric Bombieri–Lang conjecture for ramified covers of abelian varieties

### v1

@@ 3.3 A positivity lemma
Combine (3.4.16), (3.4.17), (3.4.18), (3.4.19), (3.4.20), we see that for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$,
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{1})\leq\frac{\varepsilon}{2}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$ (3.4.21)
holds for $n$ large enough.
Now we estimate the more subtle term $N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})$. Choose a positive integer $C_{3}$ such that $\widetilde{{\mathcal{R}}}_{0}\subseteq C_{3}(\widetilde{{\mathcal{R}}}_{0})_{\rm red}$. Then we have
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})\leq C_{3}N(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})$ (3.4.22)
By Corollary 3.17, we have
In fact, if $\psi_{n}(z)\notin\widetilde{{\mathcal{R}}}_{0}$, then ${\rm ord}_{z}\psi_{n}^{*}(\widetilde{{\mathcal{R}}}_{0})_{\rm red}=0\leq{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})-\min\{1,{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})\}$. If $\psi_{n}(z)\in\widetilde{{\mathcal{R}}}_{0}$, then ${\rm ord}_{z}\psi_{n}^{*}(\widetilde{{\mathcal{R}}}_{0})_{\rm red}\leq{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})$ and
$\min\{1,{\rm ord}_{z}\psi_{n}^{*}(\widetilde{{\mathcal{R}}}_{0})_{\rm red}\}=\min\{1,{\rm ord}_{z}\varphi_{n}^{*}({\mathcal{D}}_{\rm red})\}=1$. So we always have
Hence by the definition of the counting function, we get (3.4.24).
Combine (3.4.22), (3.4.23), (3.4.24), we see that
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})$ $\leq C_{3}N(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})$
$=C_{3}[N(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})-N^{(1)}(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})+N^{(1)}(r,\psi_{n},(\widetilde{{\mathcal{R}}}_{0})_{\rm red})]$
$\leq 2C_{3}(N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})).$ (3.4.25)
Recall that ${\mathcal{D}}_{t_{\infty}}^{\circ}\subseteq{\mathcal{D}}_{t_{\infty}}$ is a Zariski dense open subset of ${\mathcal{D}}_{t_{\infty}}$ satisfying the conditions in Lemma 3.7 (applying to ${\mathcal{D}}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$ and the sequence $\{t_{n}\}_{n\geq 1}$). By the property (3) of the set $S_{0}\subseteq B(\mathbb{C})$, we see that ${\mathcal{D}}_{b,t_{\infty}}^{\circ}$ is a Zariski dense open subset of ${\mathcal{D}}_{b,t_{\infty}}$.
We prove the following sub-lemma.
Sub-lemma 3.20.
Proof.
The assumption ${\rm ord}_{w_{n}}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}>1$ means that $({\rm d}\varphi^{\prime}_{n})_{0}(v_{\rm st})\in T_{\varphi^{\prime}_{n}(0)}{\mathcal{D}}_{\rm red}$ for all $n$ (large enough). By the uniform converging property, $({\rm d}\varphi^{\prime}_{n})_{0}(v_{\rm st})$ tends to $({\rm d}\varphi^{\prime}_{\infty})_{0}(v_{\rm st})$. Since $y\in{\mathcal{D}}_{b,t_{\infty}}^{\circ}$, by Lemma 3.7, we see that $({\rm d}\varphi^{\prime}_{\infty})_{0}(v_{\rm st})\in T_{\varphi^{\prime}_{\infty}(0)}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$. Or equivalently, $({\rm d}\varphi_{\infty})_{z}(v_{\rm st})\in T_{\varphi_{\infty}(z)}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$, which contradicts to the assumption that ${\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=1$.
Now take a small neighbourhood $V_{z}$ of $z$ contained in ${\mathbb{D}}_{2|z|}$ such that $\varphi_{\infty}(V_{z})\subseteq U_{y}$. Use the uniformly converge again, we see that for $n$ large enough, $\varphi_{n}(V_{z})\subseteq U_{y}$. Therefore by the argument above, ${\rm ord}_{z^{\prime}}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}\leq 1$ holds for all $z^{\prime}\in V_{z}$ and sufficiently large $n$. This finishes the proof of the sub-lemma. ∎
Back to the proof of Lemma 3.19. For fixed $r>0$ such that $\varphi_{\infty}(\partial{\mathbb{D}}_{r})\cap{\mathcal{D}}=\emptyset$ (we can enlarge $\Sigma_{\varepsilon}$ so that each $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$ satisfies this property), choose an open neighbourhood $W_{z}$ of $z$ (depending on $r$) for each $z\in\overline{{\mathbb{D}}}_{r}$ satisfying the following properties:
  1. 1.
  2. 2.
    $\log r-\log|z^{\prime}|\leq 2(\log r-\log|z|)$
  3. 3.
    If $|z|<1$, then $W_{z}\subseteq{\mathbb{D}}_{1}$.
  4. 4.
    If $\varphi_{\infty}(z)\in{\mathcal{D}}_{b,t_{\infty}}^{\circ}$ and ${\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=1$, then $W_{z}\subseteq V_{z}$, where $V_{z}$ is in sub-lemma 3.20.
By definition, we have
$\min\{\log r,\log r-\log\left|z\right|\}$ (3.4.26)
For $1\leq j\leq m$, define
By Rouché’s theorem again, for $n$ large enough, we have
Now for $n$ large enough and every $1\leq j\leq m$, if $\varphi_{\infty}(z_{j})\in{\mathcal{D}}_{b,t_{\infty}}^{\circ}$ and ${\rm ord}_{z_{j}}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=1$, we see from Sub-lemma 3.20 and assumption (4) that ${\rm ord}_{z}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}-\min\{1,{\rm ord}_{z}\varphi_{n}^{*}{\mathcal{D}}_{\rm red}\}=0$ for all $z\in W_{z_{j}}$. Thus we have $S_{j}=0$. If ${\rm ord}_{z_{j}}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=0$, then ${\rm ord}_{z_{j}}\varphi_{\infty}^{*}{\mathcal{D}}_{\rm red}=0$ so by the previous paragraph we also get $S_{j}=0$ for $n$ large enough.
for $n$ large enough.
Choose a positive integer $C_{4}$ such that $({\mathcal{D}}_{\rm red})_{b,t_{\infty}}\subseteq C_{4}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$. Then we have
${\rm ord}_{z}\varphi_{\infty}^{*}{\mathcal{D}}_{\rm red}\leq C_{4}{\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$. Moreover, if $z=z_{j}$ for some $j\in J_{2}$, then
${\rm ord}_{z}\varphi_{\infty}^{*}{\mathcal{D}}_{\rm red}\leq C_{4}{\rm ord}_{z}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red}=C_{4}\leq C_{4}{\rm ord}_{z}\varphi_{\infty}^{*}Z$. Hence by (3.4) we get
$\sum\limits_{j\in J_{1}}S_{j}+\sum\limits_{j\in J_{2}}S_{j}$ $\leq 2C_{4}\sum\limits_{j\in J_{1}}({\rm ord}_{z_{j}}\varphi_{\infty}^{*}({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\min\{\log r,\log r-\log|z_{j}|\}$
$+2C_{4}\sum\limits_{j\in J_{2}}({\rm ord}_{z_{j}}\varphi_{\infty}^{*}Z)\min\{\log r,\log r-\log|z_{j}|\}$
$\leq 2C_{4}[2(N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})-N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}))+N(r,\varphi_{\infty},Z)]$ (3.4.30)
By (3.4.7), with $\varepsilon$ replaced by $\frac{\varepsilon}{96C_{3}C_{4}}$, we get
$N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})-N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\leq\frac{\varepsilon}{32C_{3}C_{4}}N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$
holds for all $r>0$ outside a set of finite Lebesgue measure. Enlarge $\Sigma_{\varepsilon}$ if necessary, we may assume that this holds for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$.
By Theorem 3.4 we easily see that $N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\leq 2T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$ for $r$ sufficiently large. Thus we can enlarge $\Sigma_{\varepsilon}$ so that this holds for all $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$. Therefore we get
$N(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})-N^{(1)}(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})\leq\frac{\varepsilon}{16C_{3}C_{4}}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$ (3.4.31)
Denote $W=Z_{b}\cap(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$. Then by our choice of $x$ (see property (3)), $W$ has codimension at least 2 in $(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})$. By Theorem 3.6(2), we may enlarge $\Sigma_{\varepsilon}$ so that for any $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$,
$N(r,\varphi_{\infty},W)$ $\leq\frac{\varepsilon}{8C_{3}C_{4}}T(r,\varphi_{\infty},(x+{\mathcal{G}}_{b}\times\{t_{\infty}\})\cap({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$
$=\frac{\varepsilon}{8C_{3}C_{4}}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$
Thus we have $N(r,\varphi_{\infty},Z)=N(r,\varphi_{\infty},W)\leq\frac{\varepsilon}{8C_{3}C_{4}}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$. Combine this with (3.4.22), (3.4.27), (3.4.29), (3.4), (3.4.31), we get
$N(r,\psi_{n},\widetilde{{\mathcal{R}}}_{0})\leq\frac{\varepsilon}{2}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$
for any $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$ and $n$ large enough (depending on $r$). Combine with (3.4.21), we finally get
[…]
$\leq\frac{\varepsilon}{2}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})+\frac{\varepsilon}{2}T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$
$=\varepsilon T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$
for any $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$ and $n$ large enough (depending on $r$). This finishes the proof of Lemma 3.19. ∎
Now choose $0<\varepsilon<\frac{1}{2}$ such that $\frac{1}{2}(\frac{c_{1}}{2}-2\varepsilon c_{3}c_{4})>\varepsilon$. Choose $r\in\mathbb{R}^{+}\backslash\Sigma_{\varepsilon}$ large enough such that $T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})>0$. Then for $n$ large enough (depending on $\varepsilon$ and $r$) we get a contradiction from (3.4.14) and (3.4.15). This ends the proof of Proposition 3.1.
@@ 4 Proof of the main theorem
Now notice that ${\rm Sp}(X)\subsetneqq X$ by our assumption. Thus by [XY1, Corollary 4.3], which is essentially due to Ueno and Kawamata, $X$ is of general type. This means that $X$ and $\{x_{n}\}_{n\geq 1}$ satisfy the conditions in Proposition 3.1. Hence by our assumption the height of $\{x_{n}\}_{n\geq 1}$ should be bounded, a contradiction!
@@
Let $K$ be a function field of transcendence degree 1 over a field $k$ of characteristic 0 such that $k$ is algebraically closed in $K$. Let $f\colon X\to A$ be a finite map from a projective variety $X$ to an abelian variety AA ((over $K)$. Assume that $X(K)$ is Zariski dense in $X$. And there is a Zariski open dense subset $U$ of $X$ such that the heights of the points in $U(K)$ are bounded.
@@
The proof is essential the same as in [XY1, §2.2, Constancy 1 and §4.2, Constancy], we summarize it here. We first prove that $X/K$ is dominated by a constant family, and all the rational points come from this process. Namely, there is a smooth projective scheme $V$ over $k$, together with a dominant $K$-morphism
@@
is surjective. Here we denote $V_{K}=V\times_{k}K$. Let $K$ be the function field of a smooth projective curve $B$ over $k$, and let $\mathcal{X}$} be an integral model of $\mathcal{X}$} over $B$. Let $\mathcal{L}$} be an ample line bundle on $\mathcal{X}$}. Every point $x\in X(K)$ extends to a section $\tilde{x}\subseteq{\mathcal{X}}$ over $B$. By our assumption, there is a constant $c>0$ such that $\deg_{{\mathcal{L}}}(\tilde{x})<c$ for every $x\in U(K)$.
@@
contains $U(K)$. Since $U$ is Zariski dense in $X$ and $X(K)$ is Zariski dense in $X$, $U(K)$ is also Zariski dense in $X$. Thus the morphism $S\times_{k}B\to{\mathcal{X}}$ is dominant. Taking a base change by $\rm Spec$} $K\to B$, we obtain a dominant $K$-morphism $S\times_{k}K\to X$.
We can adjust the quasi-projective scheme $S$ to a smooth projective scheme over $k$. In fact, by taking a disjoint union of irreducible components, we can assume that $S$ is a disjoint union of quasi-projective varieties. Let $V^{\prime}$ be a disjoint union of projective varieties containing $S$ as a dense and open subscheme. Let $V$ be a resolution of singularities of $V^{\prime}$ by Hironaka’s theorem. The morphism $S_{K}\to X$ gives a dominant rational map $V_{K}\dashrightarrow X$. It can be extended to a morphism $V_{K}\to X$ since $X$ does not contain any rational curve, as a consequence of the finite morphism $f\colon X\to A$.
@@
Note that the normalization of $X$ also satisfies the conditions in the lemma since the normalization map is finite. Therefore to prove the constancy of the normalization of $X$, we may assume that $X$ is itself normal. In this case, the proof is the same as [XY1, §4.2, Constancy], we omit it here. This finishes the proof of the first part of the lemma. ∎
@@
Let $k_{1}$ be an intermediate field of $K/k$ such that $K/k_{1}$ has transcendence degree 1 and that $k_{1}$ is algebraically closed in $K$. As $X(K)$ is Zariski dense in $X$, it is not empty. In particular, we have an Albanese morphism $f\colon X\to{\rm Alb}(X)$ defined over $K$. Assume that Theorem 1.3 holds for the datum $(K/k_{1},X)$. Then the Albanese variety of $X$ is defined over $k_{1}$, and $X$ is birational to $(X_{1})_{K}$ for some projective variety $X_{1}$ over $k_{1}$. Moreover, the complement of the image of ${\rm Im}(X_{1}(k_{1})\to(X_{1})_{K}(K))$ in $X(K)$ (via a birational map from $(X_{1})_{K}$ to $X$) is not Zariski dense in $X$. Thus $X_{1}$ is of general type with $X(k_{1})$ Zariski dense in $X_{1}$. After taking a normalization of $X_{1}$, we can assume that $X_{1}$ is normal. Then $X_{1}$ is also of maximal Albanese dimension. In fact, after composing a birational map, the base change of the Albanese map of $X_{1}$ (over $k_{1}$) to $K$ is just the Albanese map of $X$. If in addition $f$ is finite, then there is a $K$-isomorphism $X_{1}\times_{k_{1}}K\to X$ for some projective variety $X_{1}$ over $k_{1}$. By [XY1, Lemma 4.8], the finite morphism $X_{1}\times_{k_{1}}K\to X\to{\rm Alb}(X)$ induces a finite $k_{1}$-morphism $X_{1}\to{\rm Alb}(X)^{(K/k_{1})}$. Therefore we can reduce Theorem 1.3 to the datum $(k_{1}/k,X_{1})$. Repeat the process on $k_{1}/k$, we eventually get the case ${\rm trdeg}(K/k)=1$.

### v2

@@ 3.5 Step 3: upper bound of transversal intersections
Combine (3.5.13), (3.5.14), (3.5.15), (3.5.16), (3.5.17), we see that for all $r\in\mathbb{R}_{>0}\backslash\Sigma_{\varepsilon}$, the inequality (3.5.12) holds for $n$ large enough. ∎
Proof of Lemma 3.5.1.
Recall that we have defined constants $c_{1},c_{2},c_{3},c_{4},c_{5}>0$ which are independent of $\varepsilon$. Now take $c=\frac{c_{1}}{8c_{5}}$ and choose $0<\varepsilon<\min\{\frac{1}{2},\frac{c_{1}}{2c_{3}c_{4}},\frac{c_{1}}{8(c_{3}c_{4}+1)}\}$. Take $r>1$ outside the set $\Sigma_{\varepsilon}$ such that $T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})>0$. Then we deduce from (3.5.7), (3.5.8), (3.5.12) that
$N(r,\varphi_{n},{\mathcal{D}}_{\rm red})-N^{(1)}(r,\varphi_{n},{\mathcal{D}}_{\rm red})$ $\geq\frac{1}{c_{5}}\left(\frac{c_{1}}{4}-\varepsilon(c_{3}c_{4}+1)\right)T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red})$
$\geq c\cdot T(r,\varphi_{\infty},({\mathcal{D}}_{b,t_{\infty}})_{\rm red}).$
for $n$ large enough (depending on $\varepsilon$ and $r$). This completes the proof. ∎
@@ 4 Proof of the main theorem
Now notice that ${\rm Sp}(X)\subsetneqq X$ by our assumption. Thus by [34, Corollary 4.3], which is essentially due to Ueno and Kawamata, $X$ is of general type. This means that $X$ and $\{x_{n}\}_{n\geq 1}$ satisfy the conditions in Proposition 3.1 (with $T$ taken to be trivial). Hence by our assumption the height of $\{x_{n}\}_{n\geq 1}$ should be bounded, a contradiction!
@@
Let $K$ be a function field of transcendence degree 1 over a field $k$ of characteristic 0 such that $k$ is algebraically closed in $K$. Let $X$ be a projective variety $X$ over $K$ which does not contain rational curves. Assume that $X(K)$ is Zariski dense in $X$. And there is a Zariski dense subset $U$ of $X(K)$ such that the height of the points in $U$ is bounded.
@@
The proof is essential the same as in [34, §2.2, Constancy 1 and §4.2, Constancy], we summarize it here. We first prove that $X/K$ is dominated by a constant family, and all the rational points come from this process. Namely, there is a smooth projective scheme $V$ over $k$, together with a dominant $K$-morphism
@@
is surjective. Here we denote $V_{K}=V\times_{k}K$. Let $K$ be the function field of a smooth projective curve $B$ over $k$, and let $\mathcal{X}$} be an integral model of $\mathcal{X}$} over $B$. Let $\mathcal{L}$} be an ample line bundle on $\mathcal{X}$}. Every point $x\in X(K)$ extends to a section $\tilde{x}\subseteq{\mathcal{X}}$ over $B$. By our assumption, there is a constant $c>0$ such that $\deg_{{\mathcal{L}}}(\tilde{x})<c$ for every $x\in U$.
@@
contains $U$. Since $U$ is Zariski dense in $X$, the morphism $S\times_{k}B\to{\mathcal{X}}$ is dominant. Taking a base change by ${\rm Spec}\,K\to B$, we obtain a dominant $K$-morphism $S\times_{k}K\to X$.
We can adjust the quasi-projective scheme $S$ to a smooth projective scheme over $k$. In fact, by taking a disjoint union of irreducible components, we can assume that $S$ is a disjoint union of quasi-projective varieties. Let $V^{\prime}$ be a disjoint union of projective varieties containing $S$ as a dense and open subscheme. Let $V$ be a resolution of singularities of $V^{\prime}$ by Hironaka’s theorem. The morphism $S_{K}\to X$ gives a dominant rational map $V_{K}\dashrightarrow X$. It can be extended to a morphism $V_{K}\to X$ since $X$ does not contain any rational curve.
@@
Note that the normalization of $X$ also satisfies the conditions in the lemma since the normalization map is finite. Therefore to prove the constancy of the normalization of $X$, we may assume that $X$ is itself normal. In this case, the proof is the same as [34, §4.2, Constancy], we omit it here. This finishes the proof of the first part of the lemma. ∎
Remark 4.3.
A natural question is that if the condition “$X$ does not contain rational curves” can be removed in Lemma 4.2 (with the conclusion replaced by “$X$ is birational to a constant variety $T_{K}$”). We will use the lemma only for those $X$ that admit a finite map to some abelian variety, in this case $X$ does not contain rational curves.
@@
Let $k_{1}$ be an intermediate field of $K/k$ such that $K/k_{1}$ has transcendence degree 1 and that $k_{1}$ is algebraically closed in $K$. As $X(K)$ is Zariski dense in $X$, it is not empty. In particular, we have an Albanese morphism $f\colon X\to{\rm Alb}(X)$ defined over $K$. Assume that Theorem 1.3 holds for the datum $(K/k_{1},X)$. Then the Albanese variety of $X$ is defined over $k_{1}$, and $X$ is birational to $(X_{1})_{K}$ for some projective variety $X_{1}$ over $k_{1}$. Moreover, the complement of the image of ${\rm Im}(X_{1}(k_{1})\to(X_{1})_{K}(K))$ in $X(K)$ (via a birational map from $(X_{1})_{K}$ to $X$) is not Zariski dense in $X$. Thus $X_{1}$ is of general type with $X(k_{1})$ Zariski dense in $X_{1}$. After taking a normalization of $X_{1}$, we can assume that $X_{1}$ is normal. Then $X_{1}$ is also of maximal Albanese dimension. In fact, after composing a birational map, the base change of the Albanese map of $X_{1}$ (over $k_{1}$) to $K$ is just the Albanese map of $X$. If in addition $f$ is finite, then there is a $K$-isomorphism $X_{1}\times_{k_{1}}K\to X$ for some projective variety $X_{1}$ over $k_{1}$. By [34, Lemma 4.8], the finite morphism $X_{1}\times_{k_{1}}K\to X\to{\rm Alb}(X)$ induces a finite $k_{1}$-morphism $X_{1}\to{\rm Alb}(X)^{(K/k_{1})}$. Therefore we can reduce Theorem 1.3 to the datum $(k_{1}/k,X_{1})$. Repeat the process on $k_{1}/k$, we eventually get the case ${\rm trdeg}(K/k)=1$.